OpenStudy (twizttiez):

A cone has a radius of 5 inches and a height of 18 inches. What is the exact volume of the cone? A. 150π in^3 B. 300π in^3 C. 450π in^3 D. 540π in^3

OpenStudy (twizttiez):

Thank you!

OpenStudy (anonymous):

yw apparently the formula is \[V=\frac{\pi}{3}r^2h\] plug in the numbers in to that formula

OpenStudy (twizttiez):

I got 471.24

TheSmartOne (thesmartone):

and the units which is inches^3

OpenStudy (anonymous):

your answers all have a \(\pi\) in them compute \[\frac{r^2h}{3}\] then stick a \(\pi\) next to it

TheSmartOne (thesmartone):

However, your answer choices have pi in it :P

OpenStudy (anonymous):

i.e. compute \[\frac{5^2\times 18}{3}\]

OpenStudy (anonymous):

cancel first, multiply last

OpenStudy (anonymous):

\[\frac{5^2\times 18}{3}=25\times 6=...\]

OpenStudy (twizttiez):

So A?

OpenStudy (anonymous):

yeah A

OpenStudy (anonymous):

that is rare, it is almost always C, but in this case it is A

TheSmartOne (thesmartone):

^ LOL

OpenStudy (twizttiez):

Thank you

OpenStudy (anonymous):

yw